In a reaction to determine the calcium content of hard water, a student reacts 50ml of a mono- protic acid (HCl) with 20 mls of water. What is the concentration of Calcium ions in the water if all the calcium reacts with the acid?
1
Expert's answer
2020-03-31T12:30:48-0400

Ca(HCO3)2 + 2 HCl = CaCl2 + 2H2O + 2 CO2

C(Ca2+) = 50 × (2 × CHCl / 20) = 5 × CHCl

Answer: 5CHCl.